# Topological Indices With Degree Sequence $\mathscr{D}$ of Tree

### v1

@@ Abstract.
In this paper, we refer to a asymptotic degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. The examination of topological indices on trees gives us a general overview through bounds to find the maximum and minimum bounds which reflect the maximum and minimum number of edges incident to every vertex in the graph, Albertson index known as $\sum_{uv\in E(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$, Sigma index $\sigma(G)$ among $\mathscr{D}$ of tree $T$ when $d_{n}\geqslant\dots\geqslant d_{1}$. According to the first zegrb we show for a degree sequence of order $n=4$, $\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}\left|x_{i}-x_{i+1}\right|-(b+c)-1$.
@@ 1. Introduction
Throughout this paper, let $G(E,V)$ be a simple, connected graph, denote by $uv$ an edge connecting vertices $u$ and $v$. In 1995, Molloy and Reed in [19] define asymptotic degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Zhang et al. in [17] define the set of trees $\mathcal{T}$ among degree sequence $\mathscr{D}$ denote that by $\mathcal{T}_{\mathscr{D}}$, then when $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, in [21] given $\mathcal{T}_{\mathscr{D}}$ by $\sum_{i=1}^{n}d_{i}=n-1$. The maximum degree $\Delta$ of any tree with degree sequence $\mathscr{D}$ given by Broutin et al. in [20] by $\Delta_{\mathscr{D}}=\max\{i,\mathcal{T}_{\mathscr{D}}>0\}$.
Further, let $\deg i=\deg(i)$ the degree of a vertex $i$, the amount of diverge at $u$ equals $\deg u$, which for any tree has a center chord of a maximum of one or two points. In 1971, let be refer to the maximal number of edges by $\mathcal{E}$, Wai-Kai Chen in [11] (see also [10] refer to the graph $G$ has $n$ nodes and $c$ components, then $\mathcal{E}=\frac{1}{2}(n-c)(n-c+1)$, we can make it as a general case if $G_{i},i=1,2,\ldots,c$, is the components of $G$, each having $n_{i}$ nodes, then $\mathcal{E}=\frac{1}{2}n_{i}\left(n_{i}-1\right)$. In 1997, M. O. Albertson in [12] mention to the imbalance of an edge $uv$ by $imb(uv)=\lvert d_{u}-d_{v}\rvert$, Considered graph $G$ regular if all of its vertices have the exact same degree, then irregularity measure define in [12, 16, 18] as:
@@
Dorjsembe et al. In [8] given a relationship with irregularity of Albertson index and minimum, maximum degrees $\delta,\Delta$ of graph $G$, where contribute vital roles in determining connection, shading, component incorporation, and realisability. Among this degrees, which reflect the maximum and minimum number of edges incident to every vertex in the graph, are important elements of a graph’s sophistication and performance. It is given by: $\operatorname{irr}(G)>\frac{\delta(\Delta-\delta)^{2}.|V|}{\Delta+1}$, the degree sum formula (or handshaking lemma) is know well as doubles of edges as $2|E|$. In [14, 18] define total irregularity as $\operatorname{irr}_{T}(G)=\sum_{\{u,v\}\subseteq V(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$ and shown upper bounds on the irregularity of graphs. When $G$ is a tree, Ghalavand, A. et al. In [25] proved $\operatorname{irr}_{T}(G)\leq\frac{n^{2}}{4}\operatorname{irr}(G)$ when the bound is sharp for infinitely many graphs. The first and the second Zagreb index, $M_{1}(G)$ and $M_{2}(G)$ are defined in [13, 15, 12] as:
@@
Actually an alternative expressions introduced by M. Matejí et al. in [12, 2] for the first Zagreb index as $M_{1}(G)=\sum_{u\sim v}(d_{u}+d_{v})$. The recently introduced $\sigma(G)$ irregularity index is a simple diversification of the previously established Albertson irregularity index, in [24, 23] defined as:
@@
This paper is organized as follows. In Section 1, we observe the important concepts to our work including literature view of most related papers, in Section 2 We have provided a preface through some of the important theories and properties we have utilised in understanding the work, in Section 3 we introduced the main result of our paper and the goal of this paper about Albertson index among a degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$ of tree, we computing the advanced result of this paper of Sigma index among a degree sequence of tree.
@@ 2. Preliminaries
In this section, we considered trees in graphs with fixed number of degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$. In Lemma 2.3, it is explained when a graph is connected if it contains maximal Sigma index consider $\Delta\left(G\right)=\max\left\{{{{\deg}_{G}}(v);v\in V\left(G\right)}\right\}$, that is for any vertex in graph $G$ has a maximum degree of $n-1$, as indicated by the notation $\Delta(G)=n-1$.
@@
[6] Let be $T$ a tree, a sequence of positive integers ${d_{1}},{d_{2}},{d_{3}};\quad{d_{1}}\geq{d_{2}}\geq{d_{3}}$, Albertson index defined as:
$\operatorname{irr}=\left\{\begin{aligned} &\operatorname{irr}_{max}=(d_{1}-1)^{2}+(d_{2}-1)^{2}+(d_{3}-1)(d_{3}-2)(d_{1}-d_{3})(d_{2}-d_{3})\\ &\operatorname{irr}_{min}=(d_{1}-1)^{2}+(d_{3}-1)^{2}+(d_{2}-1)(d_{2}-2)+(d_{1}-d_{3}).\end{aligned}\right.$
@@
[1, 4] If $p\geq 1$ is an integer number and a degree sequence $d=(d_{1},d_{2},\ldots,d_{n})$ where $10\leq d\leq n-$ 1, then
$\left(\sum_{i=1}^{n}d_{i}^{p}\right)^{\frac{1}{p}}\leq(n-1)^{1-\frac{1}{p}}\sum_{i=1}^{n}d_{i}^{\frac{1}{p}}$
@@
Proposition 2.2.
[3, 18]  Let be a degree sequence $d=(d_{1},\dots,d_{n})$ where $d_{1}\geq d_{2}\geq\dots\geq d_{n}$, then we have
@@
In fact, we employ this proposition in Lemma 3.1 where the degree sequence is $d=(d_{1},d_{2},d_{3},d_{4})$, and $\operatorname{irr}_{t}(G)\leq(n-2)\operatorname{irr}(G)$
@@
Let be a sequences with $n\geq 3$ and let be order as: ${d_{n}}>d_{1}>\dots>d_{2}>d_{n-1}$, the Caterpillar tree with such order has the maximum value of $\operatorname{irr}$ among all Caterpillar trees with such degrees sequence of path vertices.
@@
$\sigma(C(n,m))=\begin{cases}2m^{3},&\quad\text{if }n=2\\ 2m^{3}+m-2,&\quad\text{if }n\geqslant 2.\\ \end{cases}$
@@
For Caterpillar tree with path vertices with degrees $d_{1},d_{2},\dots,d_{n}$, we have:
$\operatorname{irr}(G)=\left({{d_{n}}-1}\right)^{2}+\left({d_{1}-1}\right)^{2}+\sum\limits_{i=2}^{n-1}{\left({{d_{i}}-1}\right)\left({{d_{i}}-2}\right)}+\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|.$
We denote by $\delta(G),\Delta(G)$ the minimum and maximum degrees in a graph $G$, denote by $d(G)$ their average degree. Let $\mathcal{S}$ be a class of graphs, then we have $\operatorname{irr}_{\max},\operatorname{irr}_{\min}$ Albertson index of a graph $G$, where:
@@
Let $G$ has the maximal Sigma index among all connected graphs with $n$ vertices and $p$ pendant vertices, where $n,p$ are positive integers such that $31\leq p\leq n-$ 3. Then: $\Delta(G)=n-1$.
Oboudi, M., R. in [9] shows if $G$ is a connected graph then we have a sequence $d_{1}>d_{2}>\dots>d_{n}$ that has $n$ degree graph and $n\geq 3$. The first and second Zagreb indices in [24] as: Two additional Zagreb-type indices are the forgotten index in [24] as:
@@
[26] Let be a complete bipartite graph is $K_{m,n}$, then sigma index of graph is $\sigma(K_{m,n})=(mn)(n-m)^{2}$.
@@
[24] For every simple graph $G,\sigma(G)$ is an even integer, so that we have:
@@
[24] For any double star graph $S_{r,k}$, the degrees of two adjacent central vertices $u,v$ defined as $d_{u}=k\geq 3,d_{v}=r\geq 1$. Then the sigma index of $S_{r,k}$ is given by
@@
Let be $\mathscr{G}(n)$ is the set of all $n$ vertex on graph $G$, $\lambda(G)$ the index of the largest eigenvalue in graph, and let be $\bar{d}(G)$ the mean of the vertex degrees, then we have:
@@
$\operatorname{irr}\left({{S_{n}}}\right)=\left({n-2}\right)\left({n-1}\right)$
@@
Let be $\mathcal{T}$ a classes of trees, then $\sigma_{max},\sigma_{min}$ in trees with $n$ vertices by:
$\sigma(\mathcal{T})=\left\{\begin{array}[]{l}{\sigma_{\max}(\mathcal{T})}=\left({n-1}\right)\left({n-2}\right){\rm{}};n\geq 3\\ {\sigma_{\min}(\mathcal{T})}=0{\rm{;n=2}}\end{array}\right.$
According to a tree with $n$ vertices has $n-1$ edges, then we have $\sigma_{\max}(T)=n-2$. Thus, the modified total Sigma as $\sigma_{t}(T)=\frac{1}{2}\sum_{(u,v)\subseteq V(T)}\left(d_{u}-d_{v}\right)^{2}$
@@ 3. Main Result
In this section, we provide the main result in this paper, we show that in Theorem 3.6 by study properties in Lemma 3.2,3.5. We mention by $C(n,m)$ caterpillar tree with $(n,m)$ vertices, and refer to $n$ main vertices, we have $X=(x_{1},x_{2},\dots,x_{n})$ where $n=|X|$, and refer to $m$ a set of leaves adjacent to main vertices. Zhang, P., Wang, X in[5] show “Caterpillar trees” are employed in chemical graph theory for describing molecular structures, and several topological indices, which include the Albertson index, have been used to study the characteristics of molecules.
3.1. Albertson Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
@@
Let be a sequences with $n\geq 3$ and let be order as: ${d_{n}}>d_{n-1}>\dots>d_{2}>d_{1}$, the Caterpillar tree with such order has the minimum value of $\operatorname{irr}$ among all Caterpillar trees with such degrees sequence of path vertices.
@@
Let be the sequence $d=(d_{1},\dots,d_{n})$ where $d_{n}>d_{1}>d_{n-2}>d_{3}\dots>d_{2}>d_{n-1}$ or If $d_{i}\leq d_{i+1}$, then we have $d_{i+1}$ as $d_{i}+k$ where $k>0$. where $\min\{d_{n},d_{1}\}=d_{n},(d_{n}-1)^{2}=d_{n}^{2}-2d_{n}+1$, thus $(d_{i})^{2}>|d_{i}-d_{i+1}|$. Therefore we have:
$\operatorname{irr}(G)$ $=(d_{n}-d_{1})+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}+2n-6$
$=-d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4.$
@@
From this cases notice $\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n-1}^{2}+d_{n}^{2}+2n-5$ is minimum according to the constant term, then $\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}$ is minimum when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. As desire. ∎
@@
Let be $T$ a tree with a degree sequence of positive integers vertices $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$ we define for numbers $a,b,c,d$ that:
$\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}\left|x_{i}-x_{i+1}\right|-(b+c)-1.$
@@
By using Albertson index we define the sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$, then by considering $M_{1}(G)=\sum_{i=1}^{n}d_{i}^{2}$ we discus according to Proposition 2.2 the following cases.
@@
As desired. ∎
@@
Let be $T$ a tree of order $n\geqslant 4$, define the sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
@@
Let be $T$ a tree with a degree sequence of positive integers $d_{i}=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{5}$ and $d_{j}=(d_{1},\dots,d_{6})$ where $d_{1}\geqslant\dots\geqslant d_{n}$ where $i,j>0$ and let be define positive numbers $a,b,c,d,e$, Albertson index among tree $T$ is:
$\operatorname{irr}(T)=\sum_{i=1}^{5}(d_{i}-1)(d_{i}-2)+\sum_{i=1}^{5}|d_{i}-d_{i+1}|+(d_{1}-1)^{2}+(d_{5}-1)^{2}.$
@@
Let be $T$ tree of order $n\geqslant 5$, a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
@@
According to Lemma 3.2, we will employed that for a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, therefore we will discus following cases by considering the constant term is $\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)$.
@@
  • In this case, the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$, then we have
@@
  • In this case, the sequence is $d=(d_{1},d_{3},d_{2},d_{4},d_{5})$, the constant term is $\lvert d_{3}-d_{2}\rvert+\lvert d_{4}-d_{5}\rvert-2$, then we have,

### v2

# Topological Indices in Trees With Degree Sequence $\mathscr{D}$

@@ Abstract.
In this paper, we denote a degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$, ordered either non-increasingly ($d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$) or non-decreasingly ($d_{1}\leqslant d_{2}\leqslant\dots\leqslant d_{n}$). The paper sets out to prove topological indices on trees with a degree sequence $\mathscr{D}$ which it is provides bounds that reflect extremal values for the number of edges incident to vertices, guided by the Albertson index (defined as $\sum_{uv\in E(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$) and the Sigma index $\sigma(G)$ for trees $T$ with degree sequence $\mathscr{D}$.
For the first Zagreb index $M_{1}(T)$, we demonstrate that for a degree sequence $\mathscr{D}$ of order $n=4$, the irregularity measure of Albertson index satisfies:
$\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}|x_{i}-x_{i+1}|-(d_{2}+d_{3})-1.$
@@ 1. Introduction
Throughout this paper, let $G=(V,E)$ be a simple, connected graph, denote by $uv$ an edge connecting vertices $u$ and $v$. In 1995, Molloy and Reed in [21] define a degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Zhang et al. in [19] define the classes of trees $\mathcal{T}$ among a non-increasing degree sequence $\mathscr{D}$, denote that by $\mathcal{T}_{\mathscr{D}}$ when $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, in [23] define $\mathcal{T}_{\mathscr{D}}$ as $\sum_{i=1}^{n}d_{i}=n-1$. The maximum degree $\Delta$ of any tree with a degree sequence $\mathscr{D}$ defined by Broutin et al. in [22] as $\Delta_{\mathscr{D}}=\max\{i,\mathcal{T}_{\mathscr{D}}>0\}$.
Further, let $\deg v=\deg(v)$ be the degree of a vertex $v\in V(G)$, the amount of diverge at $u$ equals $\deg u$, which for any tree has a centre chord of a maximum of one or two points. For a graph $G$ with $n$ vertices and $p$ pendents, Wai-Kai Chen in [13] established that the maximum number of edges is $\Delta(E)=\frac{1}{2}(n^{2}+n+p^{2}-3p)$. For each pendent $G_{i}$ (with $n_{i}$ vertices), the maximum edges within it are $\frac{1}{2}n_{i}(n_{i}-1)$ (see also [12]).
In 1997, Albertson in [14] mention to the imbalance of an edge $uv$ by $imb(uv)$ where $imb(uv)=\lvert d_{u}-d_{v}\rvert$.We considered a graph $G$ is regular if all of its vertices have the exact same degree, then the irregularity measure defined in [14, 18, 20] as:
@@
Dorjsembe et al. In [10] given the relationship between irregularity of Albertson index and minimum, maximum degrees $\delta,\Delta$ of graph $G$ (respectfully), where contribute vital roles in determining connection, shading, component incorporation, and realisability. Among this degrees, which reflect the maximum and minimum number of edges incident to every vertex in the graph, are important elements of a graph’s sophistication and performance. The relationship given by: $\operatorname{irr}(G)>\frac{\delta(\Delta-\delta)^{2}.|V|}{\Delta+1}$, the degree sum formula (or handshaking lemma) is know well as doubles of edges as $2|E|$. In [16, 20] define total irregularity as $\operatorname{irr}_{T}(G)=\sum_{\{u,v\}\subseteq V(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$ and shown upper bounds on the irregularity of graphs. When $G$ is a tree, Ghalavand, A. et al. In [26] proved $\operatorname{irr}_{T}(G)\leq\frac{n^{2}}{4}\operatorname{irr}(G)$ when the bound is sharp for infinitely many graphs. The first and the second Zagreb index, $M_{1}(G)$ and $M_{2}(G)$ are defined in [15, 17, 14] as:
@@
Actually, an alternative expressions introduced by M. Matejí et al. in [14, 3] for the first Zagreb index as $M_{1}(G)=\sum_{u\sim v}(d_{u}+d_{v})$. The recently introduced $\sigma(G)$ irregularity index is a simple diversification of the previously established Albertson irregularity index, in [24, 25] defined as:
@@
This paper is organized as follows. In Section 1, we observe the important concepts to our work including literature view of most related papers, in Section 2 we have provided a preface through some of the important theories and properties we have utilised in understanding the work, in Section 3 we introduced the main result of our paper. The goal of this paper is provide Albertson index, Sigma index among a degree sequence $\mathscr{D}=(d_{1},\dots,d_{n})$ where it is non-increasing and non-decreasing of tree $T$.
@@ 2. Preliminaries
In this section, we study trees with $\mathscr{D}=(d_{1},\dots,d_{n})$ a degree sequence where $d_{n}\geqslant\dots\geqslant d_{1}$. Lemma 2.3 establishes that a graph is connected if it attains the maximal Sigma Index. Let $\Delta(G)=\max\{\deg_{G}(v):v\in V(G)\}$ denote the maximum degree of a vertex in $G$. A graph $G$ satisfies $\Delta(G)=n-1$ if at least one vertex has degree n1n-1.
@@
[8] Let $T$ be a tree, let $\mathscr{D}=(d_{1},d_{2},d_{3})$ be a degree sequence of positive integers where $d_{1}\geqslant d_{2}\geqslant d_{3}$. The Albertson index among $\mathscr{D}$ given by:
$\operatorname{irr}(T)=\left\{\begin{aligned} &\operatorname{irr}_{max}=(d_{1}-1)^{2}+(d_{2}-1)^{2}+(d_{3}-1)(d_{3}-2)(d_{1}-d_{3})(d_{2}-d_{3})\\ &\operatorname{irr}_{min}=(d_{1}-1)^{2}+(d_{3}-1)^{2}+(d_{2}-1)(d_{2}-2)+(d_{1}-d_{3}).\end{aligned}\right$.
@@
[1, 6] Let $p$ be a prime number where $p\geq 1$, let $\mathscr{D}=(d_{1},d_{2},\ldots,d_{n})$ be a degree sequence where $0\leq d_{i}\leq n-1$ when $i\in\mathbb{N}$, then the inequality satisfying:
$\left(\sum_{i=1}^{n}d_{i}^{p}\right)^{\frac{1}{p}}\leq(n-1)^{1-\frac{1}{p}}\sum_{i=1}^{n}d_{i}^{\frac{1}{p}},$
@@
Proposition $2.$
[5, 20]  Let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, then the inequality satisfying:
@@
In fact, we employ Proposition 2.2 for establishes Lemma 3.1 where we consider $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ a degree sequence systematized this approach, by considering $\operatorname{irr}_{T}(G)\leq(n-2)\operatorname{irr}(G)$. In Theorem 2.2 we refer to caterpillar tree had defined by El-Basil in [2] and Oscar Rojo in [4].
@@
Let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence with $n\geq 3$ and let be order as: $d_{n}>d_{1}>\dots>d_{2}>d_{n-1}$, the caterpillar tree with such order has the maximum value of $\operatorname{irr}$ among all caterpillar trees with such degrees sequence of path vertices.
@@
[…]
Where $(n,m)$ is mean $n$ vertices, $m$ pendent vertices.
@@
Let $T$ be a tree, let $C(n,m)$ be a caterpillar tree with path vertices, let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence, then Albertson index among caterpillar tree given by:
$\operatorname{irr}(T)=\left({{d_{n}}-1}\right)^{2}+\left({d_{1}-1}\right)^{2}+\sum\limits_{i=2}^{n-1}{\left({{d_{i}}-1}\right)\left({{d_{i}}-2}\right)}+\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|.$
We denote by $\delta(G),\Delta(G)$ the minimum and maximum degrees in a graph $G$, denote by $d(G)$ their average degree.
Definition 1.
Let $\mathcal{S}$ be a class of graphs, then we have $\operatorname{irr}_{\max},\operatorname{irr}_{\min}$ Albertson index of a graph $G$, where:
@@
Let $G$ has the maximal Sigma index among all connected graphs with $n$ vertices and $p$ pendant vertices, where $n,p$ are positive integers such that $1\leq p\leq n-3$. Then: $\Delta(G)=n-1$.
Oboudi, M., R. in [11] shows if $G$ is a connected graph then let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence where $d_{1}>d_{2}>\dots>d_{n}$ that has $n$ degree graph and $n\geq 3$. The first and second Zagreb indices in [24] as two additional Zagreb-type indices established the forgotten index as:
@@
[27] Let $K_{m,n}$ be a complete bipartite graph, then sigma index of $K_{m,n}$ given by $\sigma(K_{m,n})=(mn)(n-m)^{2}$.
@@
[24] For every simple graph $G$, then Sigma index $\sigma(G)$ is an even integer, so that we have:
@@
[24] Let $S_{r,k}$ be a double star graph , the degrees of two adjacent central vertices $u,v$ defined as $d_{u}=k\geq 3,d_{v}=r\geq 1$. Then the sigma index of $S_{r,k}$ is given by
@@
Let $\mathscr{G}(n)$ be is the set of all $n$ vertex on graph $G$, $\lambda(G)$ the index of the largest eigenvalue in graph, and let $\bar{d}(G)$ be the mean of the vertex degrees, then we have:
@@
$\operatorname{irr}\left({{S_{n}}}\right)=\left({n-2}\right)\left({n-1}\right).$
@@
Let $\mathcal{T}$ be a classes of trees, then $\sigma_{max},\sigma_{min}$ in trees with $n$ vertices by:
$\sigma(\mathcal{T})=\left\{\begin{array}[]{l}{\sigma_{\max}(\mathcal{T})}=\left({n-1}\right)\left({n-2}\right){\rm{}};n\geq 3\\ {\sigma_{\min}(\mathcal{T})}=0{\rm{;n=2}}.\end{array}\right$.
For any tree $T$ with $n$ vertices has n1n-1 edges, then $\sigma_{\max}(T)=n-2$. Thus, the modified total Sigma in trees as $\sigma_{t}(T)=\frac{1}{2}\sum_{(u,v)\subseteq V(T)}\left(d_{u}-d_{v}\right)^{2}$.
@@ 3. Main Result
In this section, we provide the main result of this paper, we show that in Theorem 3.6 by study properties in Lemma 3.2,3.5. We have $X=(x_{1},x_{2},\dots,x_{n})$ where $n=|X|$, and refer to $m$ a set of leaves adjacent to main vertices. Zhang, P., Wang, X in[7] show “caterpillar trees” are employed in chemical graph theory for describing molecular structures, and several topological indices, which include the Albertson index, have been used to study the characteristics of molecules.
3.1. Albertson Index With Non- Increasing Degree Sequence
Let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be a degree sequence where $d_{1}\leqslant d_{2}\leqslant\dots d_{n}$, we show in Hypothesize 1 the minimum value of Albertson index among all caterpillar trees. In Lemma 3.1 we discuss a degree sequence of tree according to the first Zagreb index and in Lemma 3.2 we have term $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$. Also in both of Lemma 3.3,3.4,3.5 then that hold the main Theorem 3.6.
@@ 3.1. Albertson Index With Non- Increasing Degree Sequence
Let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence with $n\geq 3$ and let be order as: ${d_{n}}>d_{n-1}>\dots>d_{2}>d_{1}$, the caterpillar tree with such order has the minimum value of $\operatorname{irr}$ among all caterpillar trees with such degrees sequence of path vertices.
@@
Let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence where $d_{n}>d_{1}>d_{n-2}>d_{3}\dots>d_{2}>d_{n-1}$ or If $d_{i}\leq d_{i+1}$, then we have $d_{i+1}$ as $d_{i}+k$ where $k>0$. where $\min\{d_{n},d_{1}\}=d_{n},(d_{n}-1)^{2}=d_{n}^{2}-2d_{n}+1$, thus $(d_{i})^{2}>|d_{i}-d_{i+1}|$. Therefore we have:
$\operatorname{irr}(G)$ $=\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}+d_{n}-d_{1}+2n-6$
$=\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})-d_{1}+2n-4.$
@@
From this cases notice $\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n-1}^{2}+d_{n}^{2}+2n-5$ is minimum according to the constant term, then $\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}$ is minimum when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. As desire. ∎ Actually, we provide in Proposition 3.1 the inequality of maximum and minimum degree in tree (see Lemma 2.7) according to the first Zagreb index as we show that in Lemma 3.1.
Proposition 3.1.
Let $T$ be a tree of order $n=4$, let $\mathscr{D}=(d_{1},\dots,d_{4})$ be a degree sequence where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$, then we have:
$\operatorname{irr}(T)<2\sqrt{\Delta M_{1}(T)}+\delta.$
@@
Let $T$ be a tree, let $\mathscr{D}=(d_{1},\dots,d_{4})$ be a degree sequence where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$, then we have:
$\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}\left|x_{i}-x_{i+1}\right|-(d_{2}+d_{3})-1.$
@@
By using Albertson index we define the sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$, according to Proposition 3.1 we define for numbers $a,b,c,d$, then by considering $M_{1}(G)=\sum_{i=1}^{n}d_{i}^{2}$ we discus according to Proposition 2.2 the following cases.
@@
As desired. ∎ Actually, these assertions quantify how the arrangement of degree sequence $\mathscr{D}$ along the path in a caterpillar tree $C(n,m)$ where $n$ vertices and $m$ pendent vertices effects structural indices, lending credence to the notion that a given ordering produces extremal (in this case, both of Lemma 3.2,3.4) values for non-decreasing degree sequence.
@@
Let $T$ be a tree of order $n\geqslant 4$, let $\mathscr{D}=(d_{1},\dots,d_{4})$ be a non-decreasing degree sequence where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
@@
Let $T$ be a tree, let $\mathscr{D}_{i}=(d_{1},\dots,d_{5})$ be a non-increasing degree sequence where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{5}$ and let $\mathscr{D}_{j}=(d_{1},\dots,d_{6})$ be a non-increasing degree sequence where $d_{1}\geqslant\dots\geqslant d_{6}$ where $i,j>0$, Albertson index among tree $T$ is:
$\operatorname{irr}(T_{\mathscr{D}_{i}})=\sum_{i=1}^{5}(d_{i}-1)(d_{i}-2)+\sum_{i=1}^{5}|d_{i}-d_{i+1}|+(d_{1}-1)^{2}+(d_{5}-1)^{2}.$
@@
Let $T$ be tree of order $n\geqslant 5$, let $\mathscr{D}=(d_{1},\dots,d_{5})$ be a non-decreasing degree sequence where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
@@
According to Lemma 3.2, we will employed that for a degree sequence $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, therefore we will discus following cases by considering the constant term is $\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)$.
@@
  • In this case, the sequence is $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5})$, then we have
@@
  • In this case, the sequence is $\mathscr{D}=(d_{1},d_{3},d_{2},d_{4},d_{5})$, the constant term is $\lvert d_{3}-d_{2}\rvert+\lvert d_{4}-d_{5}\rvert-2$, then we have,